OpenStudy (anonymous):

Assume f(x) = 2x + 8 and g(x) = 5. What is the value of (f + g)(3)?

OpenStudy (anonymous):

Is it just f(3) = 14 + g(3) = 5, so 19?

OpenStudy (anonymous):

I am not sure how to treat g(3) when g(x) = 5 is the equation...

OpenStudy (anonymous):

you were correct it is 19!

OpenStudy (dmndlife24):

(f + g)(3) = f(3) + g(3)

OpenStudy (anonymous):

Thank you very much! :)

OpenStudy (dmndlife24):

2(3) + 8 + 5 =19 :)